OpenStudy (anonymous):

f (x) = - x ^ 2 - 3 f(-3)

OpenStudy (anonymous):

f(-3) -(-3)^2-3 = -9-3 = -12

OpenStudy (anonymous):

just plug in -3 for x

OpenStudy (anonymous):

I don't get how to figure it out

jimthompson5910 (jim_thompson5910):